OpenStudy (anonymous):

If 70 out of 100 students play tennis, and 45 of the students play chess, what is the minimum number of students that play both tennis and chess?

OpenStudy (anonymous):

15 I believe. However, not 100% sure, be sure to ask a teacher or an adult. ;)

OpenStudy (anonymous):

@Friends1 why do you think it's 15?

OpenStudy (texaschic101):

x = who both play tennis and chess tennis = 70 - x chess = 45 - x (70 - x) + x + (45 - x) = 100 70 + 45 - x = 100 115 - x = 100 -x = 100 - 115 -x = -15 x = 15 15 play both tennis -- 70 - 15 = 55 play just tennis chess -- 45 - 15 = 30 play just chess 15 + 55 + 30 = 100 answer: 15 play both tennis and chess

OpenStudy (anonymous):

Ah....I see.....Thank you so much! :)